Suppose that a car rental agency charges a fixed amount per day plus an amount per mile for renting a car. Heidi rented a car one day and paid $58 for 220 miles. On another day she rented a car from the same agency and paid $76 for 340 miles. Find the linear function f(x) (where x is in miles and f(x) is in dollars) that the agency could use to determine its daily rental charges.

Answer by dkppathak(439)

Suppose that a car rental agency charges a fixed amount per day plus an amount per mile for renting a car. Heidi rented a car one day and paid $58 for 220 miles. On another day she rented a car from the same agency and paid $76 for 340 miles. Find the linear function f(x) (where x is in miles and f(x) is in dollars) that the agency could use to determine its daily rental charges.
let fixed charge per day is Rs X and Rs Y per miles
first equations will be
X+220y = 58 (1)
smilarly
X + 340 Y=76 (2) by subtracting equations (1) and (2)
120Y = 18
Y = 18/120= $3/20 per miles
by putting the value of y in (1)
X+220 x 3/20=58
X+33=58
X = 58-33 = 25
therefore fixed charge per day will be $25 per day and charhe per mile will be $3/20
ANswer